# Finding path and cycle counting formulae in graphs with Deep Reinforcement Learning

### v1

@@ 1 Introduction
Paths and cycles are fundamental structures in graph theory, playing a crucial role in various fields such as network analysis [1], chemistry [2], computer science [3], biology [4], and social sciences [5]. Efficiently counting paths and cycles of varying lengths is essential for understanding graph connectivity and network redundancies, and is the foundation of many graph processing algorithms, including graph learning algorithms such as some recent Graph Neural Networks (GNN) [6, 7].
This problem of counting paths and cycles has been extensively studied in the literature [8, 9, 10]. Among existing approaches, matrix-based formulae such as those proposed in [11] (see equation (1) for an example) are known to be the most efficient methods for paths and cycles of lengths up to six and seven, respectively [12]. This raises a significant open question: Can a deep learning algorithm discover more efficient formulae for counting paths?
@@
Searching for formulae within a CFG corresponds to solving a combinatorial optimization problem of possibly infinite size. In recent years, Deep Reinforcement Learning (DRL) approaches have been proposed to address such problems in domains like game playing [18, 19]. A recent success of DRL has been the discovery of more efficient matrix multiplication algorithms through a Monte Carlo Tree Search (MCTS)-based approach [20]. MCTS-based RL algorithms typically consist of two phases: an acting phase and a learning phase. During the acting phase, the agent selects actions based on a heuristic that combines MCTS exploration with a deep neural network that predicts both policy and value function to guide the tree search. The network is then updated during the learning phase to reflect search trees from multiple iterations. As the objective is to discover efficient formulae and considering that MCTS aligns with CFG sentence generation process due to its tree-based structure, such a Deep MCTS approach is particularly well-suited to searching formulae within CFGs.
@@
When applied to the path/cycle counting problem, GRL not only recovers the formulae from [11] but also discovers new ones, whose computational efficiency is improved by a factor of two to six.
@@ 2.1 Path and cycle counting
As mentioned in Section 1, path/cycle counting has been extensively tackled in the literature. In the early 1970s, [8] introduced algorithms for counting cycles up to length five at the graph level. Two decades later, [9] refined these algorithms, extending cycle counting to lengths of up to seven, and conjectured that these methods can also be adapted to count cycles at the node level. Later [11] established a relationship between the counting of $l$-cycles at the edge level and the counting of $(l-1)$-paths at the edge level using a simple formula. By deriving explicit formulae for the counting of paths of length up to six at the edge level, they were able to compute the number of cycles of length up to seven. More recently, [10] rediscovered the formulae for paths of length up to four from [11]. In contrast, [12] proposed an algorithm capable of counting cycles and paths of arbitrary lengths. However, they acknowledged that their method is slower than those presented by [9] and [11]. Specifically, since the latter algorithms are based on matrix multiplication, they exhibit a computational complexity of $O\left(n^{3}\right)$, where $n$ is the number of nodes. As noted by [12], these matrix-based approaches remain the most efficient known methods for counting paths and cycles of lengths up to six and seven, respectively.
@@
A set $B$ is called a Context-Free Language (CFL) if there exists a CFG $G$ such that $B=L(G):={w\mid w\in\Sigma^{*}\text{ and }S\overset{*}{\implies}w}$.
@@
As mentioned in Section 1, it is well known that CFGs are equivalent to PDAs [21, 22, 23, 24]. Thus the following subsection is dedicated to defining PDA.
@@ 2.3 PushDown Automaton
A PushDown Automaton (PDA) is defined as a $7$-tuple $P=(Q,\Sigma,\Gamma,\delta,q_{0},Z,F)$ where $Q$ is a finite set of states, $\Sigma$ is a finite set of symbols called the input alphabet, $\Gamma$ is a finite set of symbol called the stack alphabet, $\delta$ is a finite subset of $Q\times(\Sigma\cup\{\varepsilon\})\times\Gamma\to Q\times\Gamma^{*}$, the transition relation, $q_{0}\in Q$ is the start state, $Z\in\Gamma$ is the initial stack symbol, $F\subseteq Q$ is the set of accepting states.
In the case of PDA corresponding to CFG, the input alphabet $\Sigma$ corresponds to the terminal symbol alphabet. The stack alphabet $\Gamma$ consits of $V\cup\Sigma$, which is the union of the set of variables (non-terminal symbols) and the terminal symbols. For such a PDA, there are only two states: $q_{0}$, the initial state and, $q_{1}\in F$, the accepting state. The initial stack symbol is $Z=S$, where $S$ is the start variable of the CFG. The transition relation $\delta$ consists of two types of transitions:
@@
In the same way that production rules fully defines a CFG, the transition relation $\delta$ completely specifies a PDA. For a PDA constructed from a CFG, the transposition transitions alone are sufficient to define the automaton.
A PDA generates a string by starting in the initial state $q_{0}$, with the stack initialized to $Z$ and the generated string $s$ initialized to $\varepsilon$. The PDA then processes the top symbol $t$ of the stack according to the transition relation $\delta$. If $t\in\Sigma$, a transcription occurs: $t$ is popped from the stack and appended to the output string $s$. If $t\in V=\Gamma\setminus\Sigma$, a transposition occurs: $t$ is popped from the stack, and some $v\in\{v,(q_{0},v)\in\delta(q_{0},\varepsilon,t)\}$ is pushed onto the stack. Since $v\in\Gamma^{*}$, it may consist of multiple symbols, which are pushed onto the stack in reverse order. The process continues until the stack is empty, at which point the PDA transitions to the accepting state $q_{1}$, and the generated string $s$ is a member of the language of the corresponding CFG.
@@ 3.1 From path matrix formulae to the CFG $G_{3}$
In the work of [11], all of the proposed formulae are linear combinations of terms composed of matrix multiplications and Hadamard products (denoted by $\odot$) applied exclusively on the arguments $A$, $\mathrm{I}$, $\mathrm{J}$. For example, the matrix formula in equation (1) is used to count the number of 3-paths between two nodes in the graph. The formulae for path of length four to six can be found in Appendix D of the supplementary material.
@@ 3.3 Search in $D_{3}$ through Grammar Reinforcement Learning
As stated before, the tree structure of a sentence generation within PDA (see Figure 2) aligns with MCTS algorithm. Such algorithms have been proposed and refined over the last decade to guide the search within trees with a general heuristic [25]. In this work, we propose an MCTS-based DRL algorithm, termed Grammar Reinforcement Learning (GRL) adapted to the path counting open problem, generating sets of different sentences.
@@
Each walk begins at the start state $I_{0}=\left\{Z,\cdots,Z\right\}$, whose cardinality is the number of desired sentences, and terminates when a state contains only terminal symbols. Such terminal states are sets of sentences located in leaf nodes. For each state-action pair $(I,r)$, the algorithm tracks the visit count $N(I,r)$, the empirical rule value $Q(I,r)$, and two scalars predicted by a neural network: a policy probability $\pi(I,r)$ and a value $v(I,r)$. At each intermediate state, a rule action $r$ is selected according to the following equation:
@@
After a sufficient number of MCTS, the sequences of nodes and edges from each walk are used to train the neural network. The ratio $N(I,r)/N(I)$ provides a policy derived from MCTS exploration, while $Q(I,r)$ represents the empirical expected return for the current state. The policy is learned using a Kullback–Leibler (KL) divergence loss, and the value function is trained using a mean squared error (MSE) loss. The pseudo-code for each algorithm of GRL is provided in Appendix E. Figure 3 depicts both the acting and the learning parts.
@@ 4 Gramformer
The pseudo-code for each algorithm in this framework is provided in Appendix E. Figure 5 depicts the Gramformer process for a given input.
@@
The primary objective of this experiment was to demonstrate that GRL can successfully derive the path counting formulae $P_{l}$ proposed in [11]. Specifically, for $l=2$, GRL successfully identified the formula $P_{2}$. For path lengths $l\in{3,4,5,6}$, GRL not only derived the $P_{l}$ formulae but also discovered more efficient alternatives, denoted as $P_{l}^{*}$. These new formulae significantly reduce the time complexity of $l$-path counting by factors of 2, 2.25, 4, and 6.25, respectively. The formulae for $P_{2}^{*}$ through $P_{4}^{*}$ are provided below, while those for $P_{5}^{*}$ and $P_{6}^{*}$ can be found in Appendix D.
@@
In Appendix D, we derive the cycle-counting formulae based on the work of [11], using the relation $C_{l+1}=A\odot P_{l}$. Additionally, we provide a detailed explanation of how $P_{l}^{*}$ counts $l$-path establishing a new methodology for deriving formulae.
Since $3\text{-WL}$ cannot count the $7$-paths [13], theorem 3.1 leads to the incapability for $G_{3}$ to count it either. To go beyond the 6-paths counting, a more expressive grammar is needed.
Resource consumption
@@ Appendix A CFGs and PDAs
Even if $G_{3}$ is different from the $3\text{-WL}$ CFG proposed in [17], they share the same expressive power. Indeed, in the context of this paper, we are not limited by the depth of the CFG while the goal of the grammar reduction in [17] was to keep the expressiveness of the CFG at a given depth.
@@
First of all, we have that for any matrix $N$ and vector $w$, $Nw=(N\odot\mathrm{I})w+(N\odot\mathrm{J})w$, since a sentence in the CFG (4) consists on a sum other the resulting vector, we have with the help of proposition A.1 that vectors $(N\odot\mathrm{I})w$ and $(N\odot\mathrm{J})w$ have a better separability than $Nw$. To remove $V_{c}$ variable, we first have $\mathrm{I}=\mathrm{diag}\left(\mathds{1}\right)$. Then for any matrix $N$ and vector $w$, we have that $(N\mathrm{J}\mathrm{diag}\left(w\right))\odot\mathrm{I}=\mathrm{diag}\left((N\odot\mathrm{J})w\right)$ and $(N\mathrm{diag}\left(w\right))\odot\mathrm{I}=\mathrm{diag}\left((N\odot\mathrm{I})w\right)$. The conclusion can be made by induction. We obtain $G_{3}$ as expressive as $3\text{-WL}$. ∎
@@ Appendix B On the evaluation of GRL in the context of path counting
The existence of [11] formulas, ensure that there exist a linear combination of sentences that addresses the path counting of length up to 6.
@@ Appendix D Path and cycle counting
This section contains the proof of theorem 5.1 of Section 5 and provides a detailed explanation of how $P_{l}^{*}$ counts $l$-path.
@@
The most effective explicit formula discovered to date for calculating the number of 2-paths connecting two nodes was proposed by [11], it is
@@
Following the formula for the $l$-cycle proposed in [11], we obtain for the $3$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of 3-paths connecting two nodes was proposed by [11], it is
@@
Following the formula for the $l$-cycle proposed in [11], we obtain for the $3$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of $4$-paths connecting two nodes was proposed by [11], it is
@@
Following the formula for the $l$-cycle proposed in [11], we obtain for the $5$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of $5$-paths connecting two nodes was proposed by [11], it is
@@
Following the formula for the $l$-cycle proposed in [11], we obtain for the $6$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of $6$-paths connecting two nodes was proposed by [11], it is
@@
Following the formula for the $l$-cycle proposed in [11], we obtain for the $7$-cycle the following formula
@@
Appendix E Algorithms

### v2

# Grammar Reinforcement Learning: path and cycle counting in graphs with a Context-Free Grammar and Transformer approach

@@ 1 Introduction
Paths and cycles are fundamental structures in graph theory, playing a crucial role in various fields such as network analysis [Wang et al., 2023], chemistry [Ishida et al., 2021], computer science [AbuSalim et al., 2020], biology [Bortner and Meshkat, 2022], and social sciences [Boccaletti et al., 2023]. Efficiently counting paths and cycles of varying lengths is essential for understanding graph connectivity and network redundancies, and is the foundation of many graph processing algorithms, including graph learning algorithms such as some recent Graph Neural Networks (GNN) [Bouritsas et al., 2022, Michel et al., 2023]. In particular, Bouritsas et al. [2022] demonstrated that incorporating precomputed counts of paths or cycles, either at the node level or the graph level, into the feature representation can significantly enhance the expressive power of GNNs.
This problem of counting paths and cycles has been extensively studied in the literature [Harary and Manvel, 1971, Alon et al., 1997, Jokić and Van Mieghem, 2022]. Among existing approaches, matrix-based formulae such as those proposed in Voropaev and Perepechko [2012] (see equation (1) for an example) are known to be the most efficient methods for paths and cycles of lengths up to six and seven, respectively [Giscard et al., 2019]. This raises a significant open question: Can a deep learning algorithm discover more efficient formulae for counting paths?
@@
The use of context-free grammars (CFGs) for search has gained considerable attention in the literature. In robotics, Zhao et al. [2020] demonstrated that CFGs can effectively describe robotic structures, enabling the design and adaptation of robots for diverse tasks to be formulated as a search problem within a CFG. Similarly, in the field of AutoML, the application of CFG-based search has been explored [Klinghoffer et al., 2023, Vázquez et al., 2023]. In particular, Vazquez et al. [2022] proposed a grammar specifically designed for an AutoML task, where searching within this grammar enables the discovery of the optimal model for a given task.
Searching for formulae within a CFG corresponds to solving a combinatorial optimization problem of possibly infinite size. In recent years, Deep Reinforcement Learning (DRL) approaches have been proposed to address such problems [Vinyals et al., 2015, Khalil et al., 2017, Silver et al., 2018, Hubert et al., 2021, Darvariu et al., 2024]. A recent success of DRL has been the discovery of more efficient matrix multiplication algorithms through a Monte Carlo Tree Search (MCTS)-based approach [Fawzi et al., 2022]. MCTS-based RL algorithms typically consist of two phases: an acting phase and a learning phase. During the acting phase, the agent selects actions based on a heuristic that combines MCTS exploration with a deep neural network that predicts both policy and value function to guide the tree search. The network is then updated during the learning phase to reflect search trees from multiple iterations. As the objective is to discover efficient formulae and considering that MCTS aligns with CFG sentence generation process due to its tree-based structure, such a Deep MCTS approach is particularly well-suited to searching formulae within CFGs.
@@
When applied to the path/cycle counting problem, GRL not only recovers the formulae from Voropaev and Perepechko [2012] but also discovers new ones, whose computational efficiency is improved by a factor of two to six.
@@ 2.1 Path and cycle counting
As mentioned in Section 1, path/cycle counting has been extensively tackled in the literature. In the early 1970s, Harary and Manvel [1971] introduced algorithms for counting cycles up to length five at the graph level. Two decades later, Alon et al. [1997] refined these algorithms, extending cycle counting to lengths of up to seven, and conjectured that these methods can also be adapted to count cycles at the node level. Later Voropaev and Perepechko [2012] established a relationship between the counting of $l$-cycles at the edge level and the counting of $(l-1)$-paths at the edge level using a simple formula. By deriving explicit formulae for the counting of paths of length up to six at the edge level, they were able to compute the number of cycles of length up to seven. More recently, Jokić and Van Mieghem [2022] rediscovered the formulae for paths of length up to four from Voropaev and Perepechko [2012]. In contrast, Giscard et al. [2019] proposed an algorithm capable of counting cycles and paths of arbitrary lengths. However, they acknowledged that their method is slower than those presented by Alon et al. [1997] and Voropaev and Perepechko [2012]. Specifically, since the latter algorithms are based on matrix multiplication, they exhibit a computational complexity of $O\left(n^{3}\right)$, where $n$ is the number of nodes. As noted by Giscard et al. [2019], these matrix-based approaches remain the most efficient known methods for counting paths and cycles of lengths up to six and seven, respectively.
@@
A set $B$ is called a Context-Free Language (CFL) if there exists a CFG $G$ such that $B=L(G):=\{w\mid w\in\Sigma^{*}\text{ and }S\overset{*}{\implies}w\}$.
@@
As mentioned in Section 1, it is well known that CFGs are equivalent to PDAs [Schneider, 1968, Caucal, 1995, Baeten et al., 2023, DuSell and Chiang, 2024]. Usually, PDA are language acceptor, but by relabelling the input as output of the PDA, the same PDA can be used as language generator. Thus the following subsection is dedicated to defining PDA.
@@ 2.3 PushDown Automaton
A PushDown Automaton (PDA) is defined as a $7$-tuple $P=(\mathcal{Q},\Sigma,\Gamma,\delta,q_{0},Z,F)$ where $\mathcal{Q}$ is a finite set of states, $\Sigma$ is a finite set of symbols called the input alphabet, $\Gamma$ is a finite set of symbol called the stack alphabet, $\delta$ is a finite subset of $\mathcal{Q}\times(\Sigma\cup\{\varepsilon\})\times\Gamma\to\mathcal{Q}\times\Gamma^{*}$, the transition function, $q_{0}\in\mathcal{Q}$ is the start state, $Z\in\Gamma$ is the initial stack symbol, $F\subseteq\mathcal{Q}$ is the set of accepting states.
In the case of PDA corresponding to CFG, the input alphabet $\Sigma$ corresponds to the terminal symbol alphabet. The stack alphabet $\Gamma$ consits of $V\cup\Sigma$, which is the union of the set of variables (non-terminal symbols) and the terminal symbols. For such a PDA, there are only two states: $q_{0}$, the initial state and, $q_{1}\in F$, the accepting state. The initial stack symbol is $Z=S$, where $S$ is the start variable of the CFG. The transition function $\delta$ consists of two types of transitions:
@@
In the same way that production rules fully defines a CFG, the transition function $\delta$ completely specifies a PDA. For a PDA constructed from a CFG, the transposition transitions alone are sufficient to define the automaton.
A PDA generates a string by starting in the initial state $q_{0}$, with the stack initialized to $Z$ and the generated string $s$ initialized to $\varepsilon$. The PDA then processes the top symbol $t$ of the stack according to the transition function $\delta$. If $t\in\Sigma$, a transcription occurs: $t$ is popped from the stack and appended to the output string $s$. If $t\in V=\Gamma\setminus\Sigma$, a transposition occurs: $t$ is popped from the stack, and some $v\in\{v,(q_{0},v)\in\delta(q_{0},\varepsilon,t)\}$ is pushed onto the stack. Since $v\in\Gamma^{*}$, it may consist of multiple symbols, which are pushed onto the stack in reverse order. The process continues until the stack is empty, at which point the PDA transitions to the accepting state $q_{1}$, and the generated string $s$ is a member of the language of the corresponding CFG.
@@ 3.1 From path matrix formulae to the CFG $G_{3}$
In the work of Voropaev and Perepechko [2012], all of the proposed formulae are linear combinations of terms composed of matrix multiplications and Hadamard products (denoted by $\odot$) applied exclusively on the arguments $A$, $\mathrm{I}$, $\mathrm{J}$. For example, the matrix formula in equation (1) is used to count the number of 3-paths between two nodes in the graph. The formulae for path of length four to six can be found in Appendix D of the supplementary material.
@@ 3.3 Search in $D_{3}$ through Grammar Reinforcement Learning
As stated before, the tree structure of a sentence generation within PDA (see Figure 2) aligns with MCTS algorithm. Such algorithms have been proposed and refined over the last decade to guide the search within trees with a general heuristic [Świechowski et al., 2023]. In this work, we propose an MCTS-based DRL algorithm, termed Grammar Reinforcement Learning (GRL) adapted to the path counting open problem, generating sets of different sentences.
@@
Each walk begins at the start state $I_{0}=\left\{Z,\cdots,Z\right\}$, whose cardinality is the number of desired sentences, and terminates when a state contains only terminal symbols. Such terminal states are sets of sentences located in leaf nodes. For each state-action pair $(I,r)$, the algorithm tracks the visit count $N(I,r)$, the empirical rule value $Q(I,r)$, and two scalars predicted by a neural network: a policy probability $\pi(I,r)$ and a value $v(I,r)$. From a reinforcement learning perspective, $\frac{Q(I,r)}{N(I,r)}$ represents the expected return over all possible trajectories that originate from state $I$ and follow action $r$. At each intermediate state, a rule action $r$ is selected according to the following equation:
@@
After a sufficient number of MCTS, the sequences of nodes and edges from each walk are used to train the neural network. The ratio $N(I,r)/N(I)$ provides a policy derived from MCTS exploration, while $Q(I,r)$ represents the empirical expected return for the current state. The policy is learned using a Kullback–Leibler (KL) divergence loss, and the value function is trained using a mean squared error (MSE) loss. The pseudo-code for each algorithm of GRL is provided in Appendix F. Figure 3 depicts both the acting and the learning parts.
@@ 4 Gramformer
The pseudo-code for each algorithm in this framework is provided in Appendix F. Figure 5 depicts the Gramformer process for a given input.
@@
The primary objective of this experiment was to demonstrate that GRL can successfully derive the path counting formulae $P_{l}$ proposed in Voropaev and Perepechko [2012]. Specifically, for $l=2$, GRL successfully identified the formula $P_{2}$. For path lengths $l\in{3,4,5,6}$, GRL not only derived the $P_{l}$ formulae but also discovered more efficient alternatives, denoted as $P_{l}^{*}$. These new formulae significantly reduce the time complexity of $l$-path counting by factors of 2, 2.25, 4, and 6.25, respectively. The formulae for $P_{2}^{*}$ through $P_{4}^{*}$ are provided below, while those for $P_{5}^{*}$ and $P_{6}^{*}$ can be found in Appendix D.
@@
In Appendix D, we derive the cycle-counting formulae based on the work of Voropaev and Perepechko [2012], using the relation $C_{l+1}=A\odot P_{l}$. Additionally, we provide a detailed explanation of how $P_{l}^{*}$ counts $l$-path establishing a new methodology for deriving formulae.
Since $3\text{-WL}$ cannot count the $7$-paths [Fürer, 2017], theorem 3.1 leads to the incapability for $G_{3}$ to count it either. To go beyond the 6-paths counting, a more expressive grammar is needed.
In Appendix E, we evaluate GRL on directed graphs and compare its performance to baseline algorithms. GRL stands out as the only approach capable of discovering novel matrix formulae for counting paths of lengths 4 to 5 in directed graphs.
@@ Appendix A CFGs and PDAs
Even if $G_{3}$ is different from the $3\text{-WL}$ CFG proposed in Piquenot et al. [2024], they share the same expressive power. Indeed, in the context of this paper, we are not limited by the depth of the CFG while the goal of the grammar reduction in Piquenot et al. [2024] was to keep the expressiveness of the CFG at a given depth.
@@
First of all, we have that for any matrix $N$ and vector $w$, $Nw=(N\odot\mathrm{I})w+(N\odot\mathrm{J})w$, since a sentence in the CFG (4) consists on a sum other the resulting vector, we have with the help of proposition A.1 that vectors $(N\odot\mathrm{I})w$ and $(N\odot\mathrm{J})w$ have a better separability than $Nw$. To remove $V_{c}$ variable, we first have $\mathrm{I}=\mathrm{diag}\left(\mathds{1}\right)$. Then for any matrix $N$ and vector $w$, we have that $(N\mathrm{diag}\left(w\right)\mathrm{J})\odot\mathrm{I}=\mathrm{diag}\left((N\odot\mathrm{J})w\right)$ and $(N\mathrm{diag}\left(w\right))\odot\mathrm{I}=\mathrm{diag}\left((N\odot\mathrm{I})w\right)$.
$(N\mathrm{diag}\left(w\right)\mathrm{J})_{i,i}$ $=\sum_{l,m}N_{i,l}\mathrm{diag}\left(w\right)_{l,m}\mathrm{J}_{m,i}$
$=\sum_{l}N_{i,l}w_{l}\mathrm{J}_{l_{i}}$
$=\sum_{l}N_{i,l}\mathrm{J}_{i_{l}}w_{l}$
$=\sum_{l}(N\odot\mathrm{J})_{i,l}w_{l}=((N\odot\mathrm{J})w)_{i},$
$(N\mathrm{diag}\left(w\right))_{i,i}$ $=\sum_{l}N_{i,l}\mathrm{diag}\left(w\right)_{l,i}$
$=N_{i,i}w_{i}$
$=\sum_{l}(N\odot\mathrm{I})_{i,l}w_{l}=((N\odot\mathrm{I})w)_{i}$
The conclusion can be made by induction. We obtain $G_{3}$ as expressive as $3\text{-WL}$. ∎
@@ Appendix B On the evaluation of GRL in the context of path counting
The existence of Voropaev and Perepechko [2012] formulas, ensure that there exist a linear combination of sentences that addresses the path counting of length up to 6.
@@ Appendix D Path and cycle counting
This section contains the proof of theorem 5.1 of Section 5 and provides a detailed explanation of how $P_{l}^{*}$ counts $l$-path. In the following, all graphs are assumed to be simple, i.e., they contain no self-loops. This assumption aligns with the search for paths, as self-loops cannot contribute to any path due to the repetition of a node when traversing through a self-loop. Consequently, the adjacency matrix $A$ will always have a zero diagonal. Furthermore, this ensures that $\mathrm{J}\odot A=A$.
@@
The most effective explicit formula discovered to date for calculating the number of 2-paths connecting two nodes was proposed by Voropaev and Perepechko [2012], it is
@@
Following the formula for the $l$-cycle proposed in Voropaev and Perepechko [2012], we obtain for the $3$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of 3-paths connecting two nodes was proposed by Voropaev and Perepechko [2012], it is
@@
Following the formula for the $l$-cycle proposed in Voropaev and Perepechko [2012], we obtain for the $3$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of $4$-paths connecting two nodes was proposed by Voropaev and Perepechko [2012], it is
@@
Following the formula for the $l$-cycle proposed in Voropaev and Perepechko [2012], we obtain for the $5$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of $5$-paths connecting two nodes was proposed by Voropaev and Perepechko [2012], it is
@@
Following the formula for the $l$-cycle proposed in Voropaev and Perepechko [2012], we obtain for the $6$-cycle the following formula
@@
The most effective explicit formula discovered to date for calculating the number of $6$-paths connecting two nodes was proposed by Voropaev and Perepechko [2012], it is
@@
Following the formula for the $l$-cycle proposed in Voropaev and Perepechko [2012], we obtain for the $7$-cycle the following formula
@@
We evaluated the computational time of $P_{3}^{*}$ and $P_{6}^{*}$ and the precomputation times of GSN on the IMDB-MULTI dataset, with the results shown in Figure 14. This analysis underscores the efficiency gains provided by the formulae.
Figure 14: Time computation of GRL formulae and preconsumption of GSN (log scale) on IMDB-MULTI dataset for path of length 3 and 6.
Appendix E GRL applied on directed graphs
In this section, we extend the application of GRL to derive matrix formulae for counting paths and cycles in directed graphs.
To adapt the grammar reduction approach proposed in Piquenot et al. [2024] for directed graphs, minor modifications are necessary. Specifically, the transpose operation becomes critical due to the asymmetry of the adjacency matrix in directed graphs. Despite this adjustment, the proof used to eliminate the $V_{c}$ variable remains valid. Consequently, the grammar $G_{d}$ provided to GRL for this task is defined as follows.
$E$ $\rightarrow(E\odot M)\ |\ (NE)\ |\ (EN)\ |\ E^{\mathbf{T}}\ |\ A\ |\ \mathrm{J}$ (58)
$N$ $\rightarrow(N\odot M)\ |\ (N\odot N)\ |\ \mathrm{I}$
$M$ $\rightarrow(MM)\ |\ (EE)\ |\ M^{\mathbf{T}}$
In addition to GRL, we explored search within the CFG using two alternative methods: MCTS without GramFormer, referred to as MC, and a completely random rule selection process, referred to as Rand. For both MC and Rand, we maintained the same rollout budget and maximum sentence length as those used for GRL. In the following, we denote the path matrix for directed graphs as $P_{l}^{d}$.
For $l=2$, since the task can be resolved by a single short formula, GRL, MC, and Rand all successfully identified $P_{2}^{d}$.
$P_{2}^{d}=\mathrm{J}\odot A^{2}.$
For $l=3$, when GRL was initially applied to directed graphs, we aimed to identify a combination of four sentences to construct the formula. However, GRL determined that, similar to the case of undirected graphs, only two sentences were required. This resulted in the following simplified formula.
$P_{3}^{d}=\mathrm{J}\odot(A(\mathrm{J}\odot A^{2}))-A\odot((A\odot A^{\mathbf{T}})\mathrm{J}).$
Both MC and Rand failed to converge when tasked with finding a combination of four sentences. However, after GRL revealed that only two sentences were necessary, we re-evaluated MC and Rand under this condition. In this case, both methods successfully converged and identified $P_{3}^{d}$.
For $l=4$, only GRL successfully discovered a solution.
$P_{4}^{d}$ $=\mathrm{J}\odot(A(\mathrm{J}\odot(A(\mathrm{J}\odot A^{2}))))-\mathrm{J}\odot(A(A\odot((A\odot A^{\mathbf{T}})\mathrm{J})))$
$\qquad-\mathrm{J}\odot((A\odot((A\odot A^{\mathbf{T}})\mathrm{J}))A)-A\odot((A\odot(A^{2})^{\mathbf{T}})\mathrm{J})+2*A\odot A^{\mathbf{T}}\odot A^{2}.$
For $l=5$, only GRL successfully discovered a solution.
$P_{5}^{d}$ $=\mathrm{J}\odot(A(\mathrm{J}\odot(A(\mathrm{J}\odot(A(\mathrm{J}\odot A^{2}))))))-\mathrm{J}\odot(A(\mathrm{J}\odot(A(A\odot((A\odot A^{\mathbf{T}})\mathrm{J})))))$
$\qquad-\mathrm{J}\odot(A(\mathrm{J}\odot((A\odot((A\odot A^{\mathbf{T}})\mathrm{J}))A)))-\mathrm{J}\odot(A(A\odot((A\odot(A^{2})^{\mathbf{T}})\mathrm{J})))$
$\qquad-(\mathrm{J}\odot A^{2})\odot((A^{\mathbf{T}}\odot A^{2})\mathrm{J})-A\odot((A\odot(A(\mathrm{J}\odot A^{2}))^{\mathbf{T}})\mathrm{J})$
$\qquad-((A\odot A^{\mathbf{T}})\mathrm{J})\odot(\mathrm{J}\odot(A(\mathrm{J}\odot A^{2})))-A\odot((A\odot A^{\mathbf{T}})\mathrm{J})-A\odot A^{\mathbf{T}}\odot((A\odot A^{\mathbf{T}})\mathrm{J})$
$\qquad-A^{\mathbf{T}}\odot A^{2}-3A\odot(A\odot A^{\mathbf{T}})^{2}+2A\odot A^{2}\odot+^{\mathbf{T}}A\odot A^{\mathbf{T}}\odot(A(\mathrm{J}\odot A^{2})){(A^{2})}$
$\qquad+\mathrm{J}\odot((A\odot(A^{2})^{\mathbf{T}})A)+2\mathrm{J}\odot((A\odot A^{\mathbf{T}}\odot(A^{2}))A)$
$\qquad+\mathrm{J}\odot((A\odot A^{\mathbf{T}})(\mathrm{J}\odot A^{2}))+((A\odot A^{\mathbf{T}})\mathrm{J})\odot(A\odot((A\odot A^{\mathbf{T}})\mathrm{J}))$
$\qquad+A\odot(((A\odot A^{\mathbf{T}})\odot((A\odot A^{\mathbf{T}})\mathrm{J}))\mathrm{J})+2*\mathrm{J}\odot(A(A\odot A^{\mathbf{T}}\odot A^{2})).$
To the best of our knowledge, the matrix formulae derived for counting paths of lengths 2 to 5 are novel. This highlights the capability of GRL to generate new and meaningful formulae. Furthermore, when assuming that the adjacency matrix $A$ is symmetric, the formulae discovered by GRL for directed graphs align perfectly with those identified for undirected graphs. This alignment leads us to conjecture that GRL has identified optimal formulae, at least in the undirected case.
Appendix F Algorithms